# Challenges and Requirements for Benchmarking Time Series Foundation Models

### v2

@@ B.3 Potential Overlaps of Correlated Train and Test Series in Dataset Collections
Appendix C Temporal Overlap of Correlated Series Experiment
This section provides implementation details of the experiment conducted to show the influence of the temporal overlap of correlated train and test series in a TSFM setting.
C.1 Data visualizations and statistics
@@
Figure 10 illustrates the relationship between training data and test data that overlap temporally. It is important to note that while the training and test periods overlap in time, they do not share the same data points, the overlap is purely temporal, meaning both datasets cover in a small part the same time period but contain different observations or series as can be seen in Table 6.
@@
C.2 Scaling and Model Tuning
Min-max scaling was applied separately to each dataset to account for differences in scale across datasets and ensure optimal model performance. Each scaler was fitted on the respective training data and subsequently used to transform the remaining data. For the S&P 500 dataset, the scaler was fitted explicitly on data from 2015 to the end of 2019 and then applied to the data from 2020. The Time-Series Transformer was implemented via the tsai library [^11] https://timeseriesai.github.io/tsai. The tuning was performed via Optuna [^12] https://github.com/optuna/optuna for 50 trials to minimize the MAE on the validation set. The training set consisted of the first 80% of NN5 Daily and Tourism Monthly datasets as well as January 1st, 2015 until June 30th, 2019 for no-leak and January 1st 2016 until June 30th, 2020 of the stock index data for leak. The validation set consisted of the remaining 20% of NN5 Daily and Tourism Monthly datasets, and the stock index data from July 1st, 2019 until December 31st, 2019 for no-leak and July 1st, 2020 until December 31st 2020 for leak.

### v3

@@ A.3 Potential Overlaps of Correlated Train and Test Series in Dataset Collections
Appendix B Temporal Overlap of Correlated Series Experiment
B.1 Data visualizations and statistics
@@
Figure 9 illustrates the relationship between training and test data that overlap temporally. It is important to note that although the training and test periods overlap in time, they do not share the same data points. The overlap is purely temporal, meaning that both datasets cover a small portion of the same time period but contain different observations or series, as shown in Table 3.
@@
Table 4: Per-seed S&P 500 MAE breakdown for indirect leak from other countries stock indices vs no-leak experiments. The percentage difference shows how much better the model with leakage data performs than the non-leakage model. For each seed, the better MAE is displayed in bold.
Seed MAE Before Split MAE After Split
No Leak Indirect Leak % Diff No Leak Indirect Leak % Diff
410 230.53 227.72 1.22% 157.85 199.33 -26.28%
1425 219.68 178.09 18.93% 120.77 82.53 31.67%
1680 193.52 187.19 3.27% 80.74 81.22 -0.60%
1825 247.76 156.12 36.99% 110.47 70.36 36.31%
2287 231.19 206.28 10.78% 102.42 100.54 1.84%
3658 219.23 156.91 28.43% 117.25 84.62 27.84%
4013 224.75 193.16 14.06% 114.19 97.98 14.20%
4507 206.45 213.70 -3.51% 93.10 84.01 9.77%
8936 189.40 199.38 -5.27% 103.59 153.67 -48.35%
9675 206.30 182.99 11.30% 82.35 132.44 -60.82%
Mean 216.88 190.15 11.62% 108.27 108.67 -1.44%
Std Dev 17.15 21.88 21.10 38.83
B.2 Data Preprocessing and Model Tuning
Min-max scaling was applied separately to each dataset to account for differences in scale across datasets. Each scaler was fitted on the respective training data and subsequently used to transform the remaining data. For the S&P 500 dataset, the scaler was fitted explicitly on data from 2015 to the end of 2019 and then applied to the data from 2020.
The Time-Series Transformer was implemented via the tsai library [^10] https://timeseriesai.github.io/tsai. Hyperparameter tuning was performed via Optuna [^11] https://github.com/optuna/optuna for 50 trials to minimize the MAE on the validation set. The training set consisted of the first 80% of NN5 Daily and Tourism Monthly datasets as well as January 1st, 2015 until June 30th, 2019 for no-leak and January 1st 2016 until June 30th, 2020 of the stock index data for leak. The validation set consisted of the remaining 20% of NN5 Daily and Tourism Monthly datasets, and the stock index data from July 1st, 2019 until December 31st, 2019 for no-leak and July 1st, 2020 until December 31st 2020 for leak.